19. Determine the volume occupied by 2.5 mol of a gas at 18 °C if the pressure is 81.8 kPa?



1
Expert's answer
2021-04-25T23:11:07-0400

Using ideal gas equation:

PV=nRT

here P(pressure)=81.8 kPa=81.8×10^3 Pa

moles=2.5

temperature=273.15+18=291.15K

Gas constant R=8.314m^3-Pa/K-mol

now, V=nRT/P = 8.314×2.5×291.5/81.8×10^3 ≈ 74 litre